TR=36.28q-2.50q^2  TC=40+60q-8q^2

find

a)profit function

b)q^*,p^*TC^*,AR^*,AC^*

c)nature of product

d) nature of market e)nature of return


1
Expert's answer
2021-06-25T10:56:50-0400

"a)Profitfunction=Totolrevenue\u2212Totalcost"


"=36.28q\u22122.50q^ \n2\n \u221240+60q\u22128q^\n2"


"Profitfunction=96.28q\u221210.5q ^\n2\n \u221240"

 "b) \ni)AR = 36.28 -2.5q"


"ii)AC = 40\/q+60-8q"


"iii) TR = P.Q"


"P = TR\/q"


"= \frac {36.28q-2.50q^2}{q}"


"P" "= 36.28-2.5q"


"MC = 60-16q"


"MR = 36.28 -5q"


"MC = MR"


"60-16q = 36.28-5q"


"60-36.28 = -5q-16q"


"= 23.72 = 11q"


"q = \frac{23.72}{11}"


"q = 2.16"

"iv) P = 36.28-2.5q"


"P = 36.28 -2.5(2.6)"

"P = 36.28-6.5"


"P = 29.78"